OpenStudy (ljkvndlfkj):

The endpoints of AB⎯⎯⎯⎯⎯ are A(1, 4) and B(5, -1). What are the coordinates of the point that divides AB⎯⎯⎯⎯⎯ in the ratio 2 : 3?

OpenStudy (thereisnousername):

x coordinate: 1+(2/3)*(5-1)=1+(2/3)*4=1+2=3 y coordinate: 4+(2/3)*(-1-4)=4+(2/3)*-5=4+3=7